OpenStudy (anonymous):

Find the ratio, reduced to lowest terms, of the volume of a sphere with a radius of 2 inches to the volume of a sphere with a radius of 6 inches.

OpenStudy (anonymous):

Because you are looking for a ratio, all the constants will cancel out. Now, spheric volume is given by (4/3)*pi*r^3. Hence, (4/3) and pi will be canceled out, and the ratio will be the cube of the radii.

OpenStudy (anonymous):

Answer: 2^3/6^3 = 8/216 = 2/54 = 1/27